OpenStudy (anonymous):

I just need someone to tell me how to solve this. Chase and Julia are purchasing a townhouse and finance $136,400 with a 20-year 4/1 ARM at 5.15% with a 3/11 cap structure. What will their payments be at the beginning of the fifth year assuming they are charged the maximum interest rate for that year?

OpenStudy (amistre64):

must be able to read the notation :)

OpenStudy (amistre64):

a strategy is: find the payment needed for a 20 year fixed loan; balance it out after 4 years (thats the 4/1 part, 4 years at the initial rate, adjusted at 1year intervals afterwards).

OpenStudy (amistre64):

the remaining balance is termed out for the remaining part of the 20 years: at an extra 3 points added to the interest (added each adjusted year but not to exceed a total of 11 points added overall.)

OpenStudy (amistre64):

so, any idea how to find the first set of payments? or do you have a formula for finding the balance after 4 years?

OpenStudy (anonymous):

Would you use the balance remaining of ARM formula?

OpenStudy (amistre64):

i use my own formulas so i got no idea what your material has :/ i find the payment of the 20 year fixed at the intial rate: \[Bk^{20*12}\frac{1-k}{1-k^{20*12}}=P\] for the initial balance, and the compounding factor k=1+r/12

OpenStudy (amistre64):

after words, i use that for 4 years of inital payments: to balance it out\[A=Bk^{4*12}-P\frac{1-k^{4*12}}{1-k}\] this amount is then used to assess the payments at r+3, for the remaining 20-4 years

OpenStudy (amistre64):

\[A(k_1)^{16*12}\frac{1-(k_1)}{1-(k_1)^{16*12}}=P_1\] for the compounding k=1+(r+3)/12 for the adjustment

OpenStudy (anonymous):

I'm sorry, i'm trying to understand that formula you use... My brain doesn't seem to want to plug in the numbers. :/

OpenStudy (amistre64):

P = 136400k^(240)(1-k)/(1-k^(240)), k=1+.0515/12 so out initial payments are about: 911.52 http://www.wolframalpha.com/input/?i=136400k%5E%28240%29%281-k%29%2F%281-k%5E%28240%29%29%2C+k%3D1%2B.0515%2F12 ----------------- balancing that out after 4 years gets us: A = 136400k^(48)-911.52*(1-k^(48))/(1-k), k=1+.0515/12 A = 119 058 http://www.wolframalpha.com/input/?i=136400k%5E%2848%29-911.52*%281-k%5E%2848%29%29%2F%281-k%29%2C+k%3D1%2B.0515%2F12 soo, readjusting for 16 years left at 8.15 gets us P = 119058k^(16*12)(1-k)/(1-k^(16*12)), k=1+.0815/12 so im going with something about 1111.70 starting at the beginning of the 5th year. http://www.wolframalpha.com/input/?i=119058k%5E%2816*12%29%281-k%29%2F%281-k%5E%2816*12%29%29%2C+k%3D1%2B.0815%2F12 youll of course want to verify with whatever formulas your material gives you to play with.

OpenStudy (anonymous):

OH! I see it now! Thank you so much! I'll try with my formula and see if the answer is the same.

OpenStudy (amistre64):

sometimes my formulas are too precise and can be off from the expected results so thats why i say its best to use the materials formulas ... which i can never remember lol

OpenStudy (anonymous):

Yea it's hard to remember all of them. Your formula looks simpler lol Thank you so much for you help! With my formula i actually got the same answer.

OpenStudy (amistre64):

same answers are comforting lol ...